INAF LBC+MODS queue observing, 2012 DATA SUMMARY February

Observer: K. Boutsia, M. Fumana
Telescope Operator: D. Gonzales-Huerta / G. Bechetti
Support Astronomer: D. Thompson / O. Kuhn
Instrument: LBC+MODS

FEBRUARY 18-19

Project Objects Filter ExpTime Notes
Valsecchi 1999AO10 g-sloan / r-sloan - tried to do Non-siderea tracking but it did not work - we sent suggestions to the PI how to adapt the OBs
Storm IC10VI V-bess / I-bess 10min aborted twice, first because the blue side was out of focus, then because seeing ~1.8'' - fwhmb=2.12 / fwhmr=1.75 - no useful data
Bedin M35_N2158 V-bess / i-sloan 20min first images within spec but after the 2nd image in an OB it could not hold the focus and would go out of spec - fwhmb=1.44 / fwhmr=1.67 - no useful data
Donati be27 B / V,I 12min fwhmB=1.83 / fwhmV=1.59 / fwhmI=1.96 - all data out of specs
Donati be30 B / V,I 12min fwhmB=1.41 / fwhmV=1.54 / fwhmI=1.04 - marginally within specs as far as it concerns seeing and red images elongated: average ellipR=0.13 with peaks of 0.2
Donati ngc2158 B / V,I 12min fwhmB=1.12 / fwhmV=1.37 / fwhmI=1.05 - average ellipticity on the red: 0.11 - this set could be considered within specs.
Gilli field1 rsloan / i-sloan,z-sloan 36min fwhmblue=1.47 / fwhmred=1.54 (there are 12min with fwhmred=1.1") - out of specs

FEBRUARY 19-20

Project Objects Filter ExpTime Notes
Gilli field1 r-sloan / i-sloan ~1hr all data out of specs - seeing >1.5"

FEBRUARY 20-21

Project Objects Filter ExpTime Notes
Valsecchi 1999AO10 g-sloan / r-sloan ~20min tried to do Non-siderea tracking but it ignores offsets. Try to hack ephemeris
Bedin M35_N2158 V-bess / i-sloan ~105min considered complete - average fwhm on blue=0.87" and red= 0.83" (from allseeing)
Bellazzini Sextans_A g-sloan / r-sloan ~25min check if it is complete - average fwhm on blue=0.89" and red= 0.82" (from allseeing)
Bellazzini Sextans_B g-sloan / r-sloan ~25min considered complete - average fwhm on blue=1.04" and red= 0.86" (from allseeing)
Gilli SDSS J1148+5251 r-sloan / i-sloan, z-sloan 3h considered complete - average fwhm on blue=1.0" and red: i-sloan= 0.97" / z-sloan= 0.86" (from allseeing)
PhotStd SA104 all filters    
PhotStd SA107 all filters    

FEBRUARY 21-22

Project Objects Filter ExpTime Notes
Valsecchi 1999AO10 g-sloan / r-sloan 1hr this time NS tracking seems to have worked
Storm IC10VI V-bess / I-bess 5min the airmass was low (~2.3) and the seeing was >2.0" so we aborted - after allseeing fwhm in blue=2.9" and red=2.6"
Donati ber27 B / V,I ~10min after allseeing fwhm in blue=1.53" and red: V= 1.45" / I= 1.36" - considered within specs
Donati ngc2355 B / V,I ~10min after allseeing fwhm in blue=1.65" and red: V= 1.66" / I=1.53 - marginally within specs
Donati czer24 B / V,I ~10min after allseeing fwhm in blue=1.8" and red: V= 1.7" / I= 1.5" - out of specs
Donati ber30 B / V,I ~10min after allseeing fwhm in blue=1.49" and red: V= 1.4" / I=1.22" - within specs
Donati coll74 B / V,I ~10min after allseeing fwhm in blue=1.97" and red: V= 2.06" / I= 1.6" - out of specs
Gilli SDSS J r-sloan / i-sloan 12min aborted because of bad seeing - after allseeing fwhm in blue=1.65" and red=1.4" - out of specs
Fiore Cosmos2 dual grating 3.5hr after monitoring GCS seeing measurement ~1.5hr is out of spec - see figure attached in the end of log
Donnarumma A1703m1 red grating ~50min we aborted the second exposure because we were in twilight (only 850sec saved) - GCS seeing >1.4" during this OB - out of specs
STD_MODS Feige34 dual/red grating

FEBRUARY 22-23

Project Objects Filter ExpTime Notes
Valsecchi 1999AO10 g-sloan / r-sloan 30min seeing was bad and we did not go on with the second pointing
Storm IC10VI V-bess / I-bess 5min seeing was >3.0", measured on the images, so we aborted
Nonino macsj0647 dual grating 1hr seeing was >1.5"
Fiore Cosmos2 dual grating 3hr seeing was ~1.7" but this is comparable to yesterdays data - we have 6.5hr total - completed (?)
Donnarumma A1703m1 red grating 1hr15min seeing was around ~1.2" for most of the time going up to 1.4" at the end
STD_MODS Feige34 dual/red grating good for all MODS programs

FEBRUARY 23-24

Project Objects Filter ExpTime Notes
Valsecchi 1999AO10 g-sloan / r-sloan 1hr seeing was better than yesterday and we did both pointings
Storm IC10VI V-bess / I-bess 5min seeing was ~2.0", measured on the images, so we aborted
Nonino macsj0647 dual grating 1hr seeing was ~1.3 which is comparable to the rest of the 4hrs - this mask is considered complete
Fiore Cosmos1 dual grating 0.5hr seeing was <1.3" - this mask is considered complete
Fiore Cosmos2 dual grating 1hr 0.5hr was within specs and the second 0.5hr was as yesterday - this mask is considered complete
Donnarumma A1703m1 red grating 1hr50min seeing was around 1.0" for most of the time - this mask is considered complete
Donnarumma A1703m2 red grating 1hr50min seeing was <1.3" - the last exposure finished after sky started getting brighter - sky in the spectra does look higher!!
STD_MODS Feige34 dual/red grating good for all MODS programs

-- KonstantinaBoutsia - 17 May 2012
Topic revision: r1 - 17 May 2012, KonstantinaBoutsia
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback